Gradebook ties and Loans suppose you want to have S60°,000 for retirement in 35 years. Your account earns 8% interest How much would you need to deposit in the account each
Answer #1

Desired sum = $600,000
Annual interest rate = 8%
Monthly interest rate = 0.667%
Time to retirement = 35 years or 420 months

Let monthly deposit be $x

$x*1.00667^419 + $x*1.00667^418 + $x*1.00667^417 + … + $x*1.00667 + $x = $600,000
$x * (1.00667^420 - 1) / 0.00667 = $600,000
$x * 2,296.13555 = $600,000
$x = $261.31

Monthly deposit is $261.31